OpenStudy (anonymous):

A pair of shoes has a list price of $75 and a 15% off discount rate. What was the price after the discount?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

My answer is 11.25

OpenStudy (anonymous):

Converting percentage to a decimal: 75% = 75 / 100 = 0.75 What is 75% of 15 = 75% * 15 = 0.75 * 15 = 11.25

OpenStudy (anonymous):

Good?

OpenStudy (anonymous):

If an item costs $132 after a 20% discount rate, what was the original price? @OwenOVOXO My answer is 165

OpenStudy (dayakar):

list price*(100-d)/100

OpenStudy (dayakar):

75*(100-15)/100 =75* 85/100

OpenStudy (dayakar):

63.75

OpenStudy (anonymous):

thanks

OpenStudy (dayakar):

it's ok

OpenStudy (anonymous):

Calculate the simple interest on a loan of $15,000 at 3.5% for 5 years. My answer is 2625 @dayakar

OpenStudy (dayakar):

I = P*T*R/100

OpenStudy (dayakar):

15000*3.5*5/100

OpenStudy (dayakar):

YOUR ANSWER IS RIGHT

OpenStudy (tkhunny):

Feel free to show how you obtained your answer.